OpenStudy (anonymous):

For the four (x,y) points (1,2), (0,5), (4,1) and (2,3), which point gives the maximum value of the function M = 2x + y? A. (0,5) B. (1,2) C. (2,3) D. (4,1)

OpenStudy (aum):

Substitute the x and y values of each answer choice into M = 2x + y and see which choice gives the maximum M value.

OpenStudy (anonymous):

okay I tried that

OpenStudy (aum):

When you tried the first answer choice what is the value of M you got?

OpenStudy (anonymous):

4

OpenStudy (aum):

And how about the rest?

OpenStudy (anonymous):

can u do it please my leg is in a cast and everything uggg it sux pwease im too lazy lol

OpenStudy (anonymous):

???????????? @aum

OpenStudy (aum):

M = 2x + y (1,2): M = 2(1) + 2 = 4 (0,5): M = 2(0) + 5 = 5 (4,1): M = 2(4) + 1 = 9 (2,3): M = 2(2) + 3 = 7 Choose the one that gives the highest M value.

OpenStudy (anonymous):

can I ask u one more? @aum pwease!

OpenStudy (anonymous):

In the linear programming graph, the feasible region is highlighted to show what? All possible errors B. All possible solutions C. The best solution D. The region where there are no solutions

OpenStudy (anonymous):

@aum @aum

OpenStudy (anonymous):

@aum

OpenStudy (aum):

I will let someone else handle it. You sent me 4 messages and tagged me 3 times while I was in the middle of helping someone else.